[Brahms-l] Shift report 20040108 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 08 2004 - 16:05:12 EST
------------------------------------------------------------
Date: Thu Jan  8 16:05:11 2004, Shift: 08:00-16:00
Supervisor: Oana Ristea, with: 
Log book#15 pages 26 - 26

Shift Summary:
------------------------------------------------------------
Beam from the beginning of this shift until 9.45 am. After MCR dumped the beam, they had several problems: RHIC quench interlock, cryogenic prblem, electrical problem. So, they called and ask if we want an access. During the access, FV, DB and B. Scheetz tried to resolve the mtpc2 problem (readout does not work), but until now is not fixed (see E-log).

Run 9474 - only FS@4 deg, 1/2B - 71 tr 6; run 9475 - non-TPC pedestal run.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 8 16:05:32 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 08 2004 - 16:05:46 EST